OpenStudy (anonymous):

Find the x and y intercept for 6x+5y=120

OpenStudy (anonymous):

(0, 24) (20, 0)

OpenStudy (anonymous):

(0, 24) is y intercept (20, 0) is x intercept To find the y intercept, plug in 0 for x and solve for y. To find the x intercept, plug in 0 for y and solve for x.

OpenStudy (anonymous):

When x = 0, y = 24 When y = 0, x = 20.

OpenStudy (anonymous):

5y=-6x+120 y=-6/5x+24 y intercept = 24 (0,24) x intercept=20(20,0)

OpenStudy (anonymous):

6x+5y=120 6(0)+5y=120 y=120/5 y =24 6x+5y=120 6x+5(0)=120 x 120/6 x = 20